OpenStudy (anonymous):

George is building a slide for a new park. He wants the height of the slide to be 8 feet, and he wants the horizontal distance along the ground to measure 16 feet. After building the slide, George decides the slide needs a vertical safety support and wants to use a 4-foot piece of wood to build the support. How far from the point where the base of the slide meets the ground should George place the vertical support?

OpenStudy (anonymous):

Okay, so the value of tan(O) is given by height(perpenicular dist.)/horiontal distance(base). The O can be calculated by arctan(8/16)=0.463 Now Let the vertical support be built at a distance of 'x' from the point where slide touches the ground. height of support=4 So, tan(O)=height/horizonal dist.= 4/x Therefore, x=4/tan(0) =4/tan(0.463) =8.639 m. (Answer)

OpenStudy (anonymous):